# Observational Scaling Laws and the Predictability of Language Model Performance

### v2

@@ Abstract
Understanding how language model performance varies with scale is critical to benchmark and algorithm development. Scaling laws are one approach to building this understanding, but the requirement of training models across many different scales has limited their use. We propose an alternative, observational approach that bypasses model training and instead builds scaling laws from $\sim$80 publically available models. Building a single scaling law from multiple model families is challenging due to large variations in their training compute efficiencies and capabilities. However, we show that these variations are consistent with a simple, generalized scaling law where language model performance is a function of a low-dimensional capability space, and model families only vary in their efficiency in converting training compute to capabilities. Using this approach, we show the surprising predictability of complex scaling phenomena: we show that several emergent phenomena follow a smooth, sigmoidal behavior and are predictable from small models; we show that the agent performance of models such as GPT-4 can be precisely predicted from simpler non-agentic benchmarks; and we show how to predict the impact of post-training interventions like Chain-of-Thought and Self-Consistency as language model capabilities continue to improve.
@@ 1 Introduction
Figure 1: Observational scaling laws generalize existing compute scaling laws which directly relate training compute to downstream capabilities (dashed line) by hypothesizing the existence of a low-rank space of LM capabilities that have a log-linear relationship with compute (center), and can be extracted directly from standardized LM benchmarks (left). This enables us to get low-cost, high-resolution scaling predictions of LMs’ complex downstream capabilities from their observable standard benchmark metrics using nearly 80 publicly accessible LMs (left to right).
@@ 3.1 Generalizing Compute Scaling Laws
Several studies [e.g., 24, 63] have generalized the functional form to analyze the scaling of LMs’ downstream performance. Specifically, let $E_{m}$ represent the normalized downstream errors of models within the range $[0,1]$, they observed a sigmoidal relationship between $\log(C_{m})$ and $E_{m}$ and thus used a logistic link function instead of a logarithm for the generalized linear model in Eq. 1:
@@ 4 Validating Observational Scaling Laws
To ensure that our scaling laws are actually predictive and that we are not simply overfitting through various choices in scaling law construction and hyperparameters, we design our experiments to have systematic holdout sets and robustness checks. We also preregister our predictions for future models after the release of the paper as a test of whether our scaling laws overfit current models. We release our code including the implementation and collected data at https://github.com/ryoungj/ObsScaling.
@@
In Sec. C.7, we include all functional forms for the fitted scaling laws in our experiments as preregistration of our predictions for future models. We will assess the accuracy of these scaling laws (without refitting) using models developed after May 2024 and commit to updating the manuscript on ArXiv with our prediction results after 4 months.
@@ 4.1 Predictability of “Emergent” Capabilities
Figure 4: “Emergent” capabilities of LMs can be accurately predicted from weaker models to stronger ones with observational scaling laws, and using PC measures as the predictor provides much more accurate predictions than using compute scale measures like training FLOPs and model size (see Fig. C.10). Two non-arithmetic and two arithmetic tasks from BigBench [75], which are identified as “emergent” in [90], are used for illustration.
@@ 4.2 Predictability of Agentic Capabilities
Figure 5: (a)-(b) The agentic capabilities of instruction-tuned LMs measured by agent benchmarks can be accurately predicted from weaker models (sub GPT-3.5) to stronger ones (e.g., GPT-4) by their PC measures. (c) The fitted weights on both benchmarks demonstrate the importance of programming capabilities (HumanEval) for the agentic capabilities of LMs.
@@ 4.3 Predicting the Impact of Post-Training Techniques
Figure 6: (a) The LM performance with and without techniques like CoT and Self-Consistency can be accurately predicted with observational scaling laws. The fitted scaling curves indicate that CoT has a better scaling behavior than SC. See Fig. C.13 for detailed per-method scaling plots and comparison with compute baselines. (b) The fitted weights demonstrate a very different pattern when CoT is applied, emphasizing general knowledge (MMLU) and programming capabilities (Humaneval).
@@ 6 Discussion and Other Applications of Observational Scaling
Our work validates the hypothesis that there is a low-dimensional space of LM capabilities that captures their scaling behaviors and can be measured via a low-rank decomposition of existing LM benchmarks. While the majority of our work focuses on applications to scaling laws and predictions, we also find that the shared, low-dimensional capabilities could potentially be used as an evaluation metric and optimization target for LMs. We discuss some of these possibilities here.
@@ Acknowledgements
We thank Zitong Yang for his assistance with an early experiment of the project. We also thank Jimmy Ba, Yann Dubois, Pavan Kapanipathi, Lisa Li, Karthik Narasimhan, Ethan Perez, Chenglei Si, Tristan Thrush, Zitong Yang, Shunyu Yao, and the Hashimoto Group for their helpful discussions or feedback on the paper draft. This project is not possible without the open-source contributions including HuggingFace, EleutherAI LM Eval Harness [27], Open LLM Leaderboard [9], EvalPlus [52], vLLM [43], LMSys Chatbot Arena Leaderboard [18], and AlpacaEval Leaderboard [47].
@@ B.1.1 Pretrained Base Models
We collected a broad set of representative open LMs covering 21 model families and a total of 77 models. These model families include Llama-2 [84], Llama [83], Llama-3 [2], Qwen1.5 [80], Qwen [8], Mistral [39], Mixtral [40], Yi [98], Gemma [79], Falcon [3], Phi [48], Pythia [11], BLOOM [92], GPT-Neo/J [12], OPT [100], MPT [81], XGLM [51], CodeLlama [71], StarCoder [46], StarCoder2 [55], DeepSeek-Coder [30]. For each model, we collected their available metadata including the number of model parameters $N$ and the amount of pretraining tokens $D$ by analyzing papers and other public information. We then estimated the training FLOPs $C$ using the simple estimate of $C\approx 6ND$ [42] for each model. Note that for models that were continually pretrained on additional data such as CodeLlama, we used the sum of the pretraining tokens and the additional continual pretraining tokens to estimate $D$. See table B.1 for the collected metadata of these models.
@@ B.1.2 Instruction-Tuned Models
Table B.1: Collected metadata and base evaluation metrics for base pretrained models used in Sec. 4.1, Sec. 4.3, and Sec. 5. Model names follow the HuggingFace naming. See data collection details in Sec. B.1.1.
@@ C.2 Properties of PC measures
C.3 Robustness Checks
@@ C.4 Emergent Capabilities
C.5 Post-Training Method Analysis
@@
C.6 Model Subset Selection
@@
C.7 Fited Functional Forms for Preregistration of Predictions
In table C.1, we included the functional forms of fitted scaling laws included in our experiments. These functional forms serve as a preregistration of our predictions for future models, which will be used to test the generalizability of our scaling analysis to unseen models.

### v3

@@ Abstract
Understanding how language model performance varies with scale is critical to benchmark and algorithm development. Scaling laws are one approach to building this understanding, but the requirement of training models across many different scales has limited their use. We propose an alternative, observational approach that bypasses model training and instead builds scaling laws from $\sim$100 publically available models. Building a single scaling law from multiple model families is challenging due to large variations in their training compute efficiencies and capabilities. However, we show that these variations are consistent with a simple, generalized scaling law where language model performance is a function of a low-dimensional capability space, and model families only vary in their efficiency in converting training compute to capabilities. Using this approach, we show the surprising predictability of complex scaling phenomena: we show that several emergent phenomena follow a smooth, sigmoidal behavior and are predictable from small models; we show that the agent performance of models such as GPT-4 can be precisely predicted from simpler non-agentic benchmarks; and we show how to predict the impact of post-training interventions like Chain-of-Thought and Self-Consistency as language model capabilities continue to improve.
@@ 1 Introduction
Figure 1: Observational scaling laws generalize existing compute scaling laws which directly relate training compute to downstream capabilities (dashed line) by hypothesizing the existence of a low-rank space of LM capabilities that have a log-linear relationship with compute (center), and can be extracted directly from standardized LM benchmarks (left). This enables us to get low-cost, high-resolution scaling predictions of LMs’ complex downstream capabilities from their observable standard benchmark metrics using nearly 100 publicly accessible LMs (left to right).
@@ 3.1 Generalizing Compute Scaling Laws
Several studies [e.g., 25, 67] have generalized the functional form to analyze the scaling of LMs’ downstream performance. Specifically, let $E_{m}$ represent the normalized downstream errors of models within the range $[0,1]$, they observed a sigmoidal relationship between $\log(C_{m})$ and $E_{m}$ and thus used a logistic link function instead of a logarithm for the generalized linear model in Eq. 1:
@@ 4 Validating Observational Scaling Laws
To ensure that our scaling laws are actually predictive and that we are not simply overfitting through various choices in scaling law construction and hyperparameters, we design our experiments to have systematic holdout sets and robustness checks. We have also preregistered our predictions for future models after the initial release of the paper as a test of whether our scaling laws overfit current models. We release our code including the implementation and collected data at https://github.com/ryoungj/ObsScaling.
@@
In the initial release of our paper (May 2024), we have preregistered our scaling predictions for future models (see preregistered functional forms in Sec. C.9) and committed to updating the manuscript on ArXiv with our prediction results after 4 months. We have assessed these predictions on new models released after the initial paper release, collected as of September 1st 2024, including most capable open models to date such as Llama 3.1-405B [24] and Qwen2-72B [105] (see the full collected model list in Sec. B.1.1), resulting in an additional test set of 20 models for robustness checks. The results are included in Fig. 4, and additional results on other tasks and new benchmarks are included in Sec. C.3.
@@ 4.1 Predictability of “Emergent” Capabilities
Figure 4: “Emergent” capabilities of LMs can be accurately predicted from weaker models to stronger ones with observational scaling laws, and using PC measures as the predictor provides much more accurate predictions than using compute scale measures like training FLOPs and model size (see Fig. C.12). Our preregistered predictions also accurately extrapolate to new models released after the initial paper release, including Llama-3.1-405B [24]. Two non-arithmetic and two arithmetic tasks from BigBench [82], which are identified as “emergent” in [98], are used for illustration.
@@ 4.2 Predictability of Agentic Capabilities
Figure 5: (a)-(b) The agentic capabilities of instruction-tuned LMs measured by agent benchmarks can be accurately predicted from weaker models (sub GPT-3.5) to stronger ones (e.g., GPT-4) by their PC measures. (c) The fitted weights ($\beta^{\top}\gamma$) on both benchmarks demonstrate the importance of programming capabilities (HumanEval) for the agentic capabilities of LMs.
@@ 4.3 Predicting the Impact of Post-Training Techniques
Figure 6: (a) The LM performance with and without techniques like CoT and Self-Consistency can be accurately predicted with observational scaling laws. The fitted scaling curves indicate that CoT has a better scaling behavior than SC. See Fig. C.15 for detailed per-method scaling plots and comparison with compute baselines. (b) The fitted weights ($\beta^{\top}\gamma$) demonstrate a very different pattern when CoT is applied, emphasizing general knowledge (MMLU) and programming capabilities (Humaneval).
@@ 6 Discussion and Other Applications of Observational Scaling
Our work validates the hypothesis that there is a low-dimensional space of LM capabilities that captures their scaling behaviors and can be measured via a low-rank decomposition of existing LM benchmarks – which interestingly connects to the item response theory in psychometrics [58] that models humans’ test performance by their fundamental abilities such as general intelligence. While the majority of our work focuses on applications to scaling laws and predictions, we also find that the shared, low-dimensional capabilities could potentially be used as an evaluation metric and optimization target for LMs. We discuss some of these possibilities here.
@@ 7 Conclusion, Limitations, and Future Work
Limitations and future work
Finally, we discuss some limitations of our approach and findings: Firstly, observational scaling laws are primarily applicable to post-training scaling analyses and do not directly translate to pretraining scenarios in the same way as standard compute-based scaling laws. Secondly, our study mostly focuses on the scaling behavior of model capabilities measured through few-shot prompting or basic prompting techniques (such as CoT, self-consistency, or simple agent scaffolding). Extending our approach to other post-training setups, including scenarios involving fine-tuning or more intensive inference-time computation [12, 81], would be valuable. Thirdly, while we have demonstrated that our observational scaling analyses can provide meaningful insights into improving particular models’ complex capabilities, a promising direction for future work would be to apply the findings from our approach, such as by deriving surrogate measures for model complex capabilities that can be used to optimize models directly and efficiently. Lastly, our assumptions do not account for potential benchmark contamination (where particular benchmark data leaks into model training) or the heterogeneity within model families (where models within the same family may have varying compute efficiencies and scaling behaviors). Investigating the impact of these assumptions on our approach would be an interesting avenue for future research.
@@ Acknowledgements
We thank Zitong Yang for his assistance with an early experiment of the project. We also thank Jimmy Ba, Yann Dubois, Honghua Dong, Pavan Kapanipathi, Lisa Li, Karthik Narasimhan, Ethan Perez, Chenglei Si, Tristan Thrush, Zitong Yang, Shunyu Yao, the Hashimoto Group, and anonymous reviewers for their helpful discussions or feedback on the paper draft. This project is not possible without the open-source contributions including HuggingFace, EleutherAI LM Eval Harness [28], Open LLM Leaderboard [8], EvalPlus [55], vLLM [45], LMSys Chatbot Arena Leaderboard [18], and AlpacaEval Leaderboard [49].
@@
09/25/2024
  • Updated preregistered prediction results with new models released after the initial paper release.
  • Added discussion on the limitations and future work of observational scaling laws.
  • Added additional analysis on our apporach in Sec. C.8.
  • Incorporated the feedback from anonymous reviewers.
@@ B.1.1 Pretrained Base Models
We collected a broad set of representative open LMs covering 21 model families and a total of 77 models. These model families include Llama-2 [92], Llama [91], Llama-3 [24], Qwen1.5 [88], Qwen [7], Mistral [41], Mixtral [42], Yi [107], Gemma [86], Falcon [2], Phi [50], Pythia [10], BLOOM [100], GPT-Neo/J [11], OPT [109], MPT [89], XGLM [54], CodeLlama [76], StarCoder [48], StarCoder2 [59], DeepSeek-Coder [31]. For preregistration test, we have collected an additional set of 20 models covering 8 families released after May 2024 and as of September 1st 2024, including Llama-3.1 [24], Qwen2 [105], DeepSeek V2 [80], Gemma-2, [87], Jamba [52], Yi-1.5 [107], etc. For each model, we collected their available metadata including the number of model parameters $N$ and the amount of pretraining tokens $D$ by analyzing papers and other public information. We then estimated the training FLOPs $C$ using the simple estimate of $C\approx 6ND$ [44] for each model. Note that for models that were continually pretrained on additional data such as CodeLlama, we used the sum of the pretraining tokens and the additional continual pretraining tokens to estimate $D$. See table B.1 for the collected metadata of these models.
@@ B.1.2 Instruction-Tuned Models
Table B.1: Collected metadata and base evaluation metrics for base pretrained models used in Sec. 4.1, Sec. 4.3, and Sec. 5. Model names follow the HuggingFace naming. See data collection details in Sec. B.1.1. For the most up-to-date results, please refer to https://github.com/ryoungj/ObsScaling/blob/main/eval_results/base_llm_benchmark_eval.csv.
@@ C.2 Properties of PC measures
C.3 Additional Preregisteration Results
Preregistered predictions on post-training analysis tasks
In Fig. C.5, we tested our preregistered predictions on the post-training analysis tasks. We observe reasonable forecasts on new models, and the predictions using PC measures outperform the ones using compute measures like training FLOPs.
(a) Trainig FLOP based scaling laws
(b) Observational scaling laws
Figure C.5: Our preregistered predictions of observational scaling laws using PC measures (with # = 3) provides reasonable forecasts for new models that are released after our initial paper release. The predictions on naive prompting is a bit off, but still align with the general trend and perform better than using compute measures like training FLOPs.
Preregistered predictions on Open LLM leaderboard v2 benchmarks
Besides testing new models on exsiting benchmarks with preregistred predictions, we also tested observational scaling laws on the new, more challenging benchmarks being used in Open LLM Leaderboard v2. In particular, we selected a subset of new tasks where at least some exsiting open models demonstrate non-trivial performance (which will exclude benchmarks like IFEval and MUSR) and where scaling predictions from base benchmarks are non-trivial (which will exclude MMLU Pro), including GPQA [74], MATH [33], and BBH [83]. We fit both observational scaling laws and compute-based scaling laws on these benchmarks and compared their extrapolation performance. Since the tasks are more challenging, it requires a larger cutoff threshold to include more data points with non-trivial performance on these tasks. We set the FLOPs cutoff to be $16.8$, $25.2$, $8.4\times 10^{21}$ for GPQA, MATH, and BBH, respectively. The results are in Fig. C.6. We find that observational scaling laws provides reasonable forecasts on these new, challenging benchmarks and outperform compute-based scaling laws when extrapolating to larger models.
(a) Training FLOP based scaling laws
(b) Observational scaling laws
Figure C.6: Observational scaling laws also provide reasonable forcasts on new, more challenging benchmarks being used in Open LLM Leaderboard v2 and outperform compute-based scaling laws.
C.4 Robustness Checks
@@ C.5 Emergent Capabilities
C.6 Post-Training Method Analysis
@@
C.7 Model Subset Selection
@@
C.8 Additional Analysis
We have received valuable feedback from anonymous reviewers and have conducted extrnsive additional analysis to address their remaining questions.
Extracting PC measures with non-matrix factorization
We note that the benchmark coefficients on our principal capability measures are not guaranteed to be non-negative, which may hinder the interpretability of the extracted components. Therefore, we conduct an additional analysis with non-negative matrix factorization (NMF) to ensure the non-negativity of the component-wise benchmark coefficients that may provide more interpretable capability dimensions. The results are included in Fig. C.19. We observed the NMF components do generally demonstrate a interpretable decomposition, as well as a positive and smooth scaling with training FLOPs within each model family (as our PC measures).
While NMF offers enhanced interpretability and positive scaling properties compared to PCA, it also has notable limitations. Firstly, unlike PCA, NMF does not enforce orthogonality among its extracted components, as evident in the observed correlation between Components 3 and 4. Consequently, the coefficients assigned to each model across dimensions may not serve as independent measures of specific capabilities. Secondly, the ordering of NMF components lacks uniqueness and intrinsic physical meaning. This contrasts with PCA components, which are systematically ordered by their explained variances. The PCA approach provides an ‘importance’ measure for each dimension and allows for controlled trade-offs between representativeness and noise inclusion by adjusting the number of PCs used in the analysis.
(a) NMF component weights
(b) NMF compoenent scaling
Figure C.19: Extracting PC measures with non-matrix factorization: More interpretable principal capability measures can be obtained by non-negative matrix factorization (NMF). (a) NMF ensures the non-negativity of the component-wise benchmark coefficients and provides an interpretable decomposition. For example, we may view component 1 and 4 as reasoning and language understanding capabilities, respectively. (b) The NMF components generally demonstrate a smooth, positive scaling with increasing FLOPs. The results also hold across other model families and components.
Figure C.20: Pushing the limit of cutoff point: The cutoff can be further pushed back on each individual task while still providing reasonable predictions. (a) Emergent capability tasks: We include three representative tasks, and the task-specific FLOPs cutoff are 25, 84, and 8 ×$\times 10^{21}$ respectively (from left to right), compared to the unified $84\times 10^{21}$ in our current setup. We also test the newly released Llama-3.1 405B (FP8) to assess the generalization to a larger scale. (b) Agentic tasks: We test on AgentBench that has more available data points with an 80/20 train/test split. The extrapolations underestimate performance to some extent, but still align with the overall observed trend.
(a) Qwen1.5
(b) OPT
Figure C.21: Scaling predictions with single-family models: For scaling prediction from FLOPs within a single family, at least 5 models are typically required for accurate extrapolation, but the performance is highly dependent on the specific setup. We test Qwen1.5 on non-algorithmic and OPT on arithmetic tasks. Both model families demonstrate accurate extrapolation on one task but not the other.
(a) Emergent capability
(b) Post-training
Figure C.22: Confidence intervals of scaling predictions: We calculate 95% confidence intervals for predictions from the non-linear regression models at each data point, and the observed data points fall within these confidence intervals. When extrapolating from very few data points above the random-guess level (e.g., in Persian QA), the confidence intervals may be wider. We include representative tasks for both emergent capability (left) and post-training analysis (right) setups.
C.9 Fitted Functional Forms for Preregistration of Predictions
In table C.1, we included the functional forms of fitted scaling laws in our experiments. These functional forms served as a preregistration of our predictions for future models at the time of the initial paper release, which has been used to test the generalizability of our scaling analysis to unseen models.